A rectangle is 7 cm wide and 0 cm long. Find te length of a diagonal of the rectangle.
Answer by wilft1(217)

what were the dimensions again? you cant have a rectangle that has a dimension of 0
if it was 7 cm wide and 10 cm long, you can find the length of the diagonal using pathagorias theorm, a^2+b^2=c^2
side A is 7 cm, side b is 10 cm long, so square both numbers to get...
49 + 100 = 149
we now know that c^2 is 149, just use a calculator to get the square root of 149 which is 12.20655562
and thats the length of the diagonal